# Preconditioning FEM discretisations of the high-frequency Maxwell equations by either perturbing the coefficients or adding absorption

### v1

@@ 6.2 Norm of inverse solution operator under perturbation
$I-{\mathcal{A}}_{2}^{-1}{\mathcal{A}}_{1}={\mathcal{A}}_{2}^{-1}({\mathcal{A}}_{2}-{\mathcal{A}}_{1})\quad\text{ and }\quad I-{\mathcal{A}}_{1}{\mathcal{A}}_{2}^{-1}=({\mathcal{A}}_{2}-{\mathcal{A}}_{1}){\mathcal{A}}_{2}^{-1}.$
@@
6.4 Proofs of Lemma 4.2 and 5.9 (about equivalence of the norms of the continuous and discrete inverses)
@@
We highlight that this argument goes back to at least [1, Lemma 3.4], but since it is short, and crucial to the proof of Lemma 5.9, we include it for completeness. The first inequality in (4.2) is an immediate consequence of the fact that $\|\cdot\|_{{\cal H}^{*}}\leq\|\cdot\|_{{\cal H}_{0}}$ (which follows from the definition of $\|\cdot\|_{{\cal H}^{*}}$ and $\|\cdot\|_{{\cal H}_{0}}\leq\|\cdot\|_{{\cal H}}$). Let $a^{+}_{\ell}(u,v):=a_{\ell}(u,v)+C_{\rm G2,\ell}(u,v)_{{\cal H}_{0}}$ and observe that $a^{+}_{\ell}(\cdot,\cdot)$ is coercive by (2.2). To prove the second inequality in (4.2) it is sufficient to prove that, given $F\in{\cal H}^{*}$, the solution of $a_{\ell}(u,v)=F(v)$ for all $v\in{\cal H}$ satisfies
@@ Acknowledgements
The author thanks Qiya Hu (Academy of Mathematics and Systems Science, Chinese Academy of Sciences) for asking him the question of whether the results of [16, 20] extend to the time-harmonic Maxwell equations, and Théophile Chaumont-Frelet (INRIA, Lille) for discussions about the discrete inf-sup constant of the Helmholtz sesquilinear form.

### v2

@@ 5.2 Norm of solution operator under perturbation
[…]
@@
5.4 Proofs of Lemma 3.2 and 4.8 (about equivalence of the norms of the continuous and discrete inverses)
@@
We highlight that this argument goes back to at least [2, Lemma 3.4], but since it is short, and crucial to the proof of Lemma 4.8, we include it for completeness. The first inequality in (3.2) is an immediate consequence of the fact that $\|\cdot\|_{{\cal H}^{*}}\leq\|\cdot\|_{{\cal H}_{0}}$ (which follows from the definition of $\|\cdot\|_{{\cal H}^{*}}$ and $\|\cdot\|_{{\cal H}_{0}}\leq\|\cdot\|_{{\cal H}}$). Let $a^{+}_{\ell}(u,v):=a_{\ell}(u,v)+C_{\rm G2,\ell}(u,v)_{{\cal H}_{0}}$ and observe that $a^{+}_{\ell}(\cdot,\cdot)$ is coercive by (2.2). To prove the second inequality in (3.2) it is sufficient to prove that, given $F\in{\cal H}^{*}$, the solution of $a_{\ell}(u,v)=F(v)$ for all $v\in{\cal H}$ satisfies
@@ Acknowledgements
The author thanks Qiya Hu (Academy of Mathematics and Systems Science, Chinese Academy of Sciences) for asking him the question of whether the results of [18, 22] extend to the time-harmonic Maxwell equations, Théophile Chaumont-Frelet (INRIA, Lille) for discussions about the discrete inf-sup constant of the Helmholtz sesquilinear form, and the referees for many useful comments.